Encyclopedia Astronautica
2008.01.06 - ISS On-Orbit Status 01/06/08


Sunday -- off-duty day for CDR Whitson, FE-1 Malenchenko and FE-2 Tani (congratulations, Dan, to the nice write-up about you and your mom in today's Washington Post!).

Ahead: Week 12 of Increment 16. Also:

Christmas Eve for tomorrow's Russian Orthodox Christmas.

CDR Whitson & FE-2 Tani started out with the daily reading of SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight) experiment data accumulated during the night, for logging and filling in questionnaire entries in the SLEEP session file on the HRF-1 laptop for downlink. (To monitor the crewmembers' sleep/wake patterns and light exposure, Dan and Peggy wear a special Actiwatch device which measures the light levels encountered by him as well as his patterns of sleep and activity throughout the Expedition. The log entries are done within 15 minutes of final awakening for seven consecutive days, as part of the crew's discretionary 'job jar' task list.)

Upon wakeup, Dan Tani also performed the last sampling of his session with the NASA/JSC experiment NUTRITION w/Repository, collecting a final urine sample for storage in the MELFI (Minus-Eighty Laboratory Freezer for ISS), Dewar 2, Tray A/2-3. The sampling kit was then stowed away. (The current NUTRITION project is the most comprehensive in-flight study done by NASA to date of human physiologic changes during long-duration space flight. It includes measures of bone metabolism, oxidative damage, nutritional assessments, and hormonal changes, expanding the previous Clinical Nutritional Assessment profile (MR016L) testing in three ways: Addition of in-flight blood & urine collection (made possible by MELFI), normative markers of nutritional assessment, and a return session plus 30-day (R+30) session to allow evaluation of post-flight nutrition and implications for rehabilitation.)

After Houston Flight Controllers started deactivating the CDRA (Carbon Dioxide Removal Assembly) early this morning (1:40-6:40am) and cooling was no longer required, the FE-2 disconnected the ITCS LTL QD (Internal Thermal Control System/Low Temperature Loop/Quick Disconnect) jumper to the CDRA rack (LAB1D6).

With the Russian O2 (oxygen) generator 'Elektron' turned off since 12/28/07, Yuri had time set aside for another 1-hour O2 repress of the cabin atmosphere from Progress M-62/27P storage tankage, if required. (The Elektron will remain powered down until 1/9/08 to conserve hardware lifetime. During this time, the station will be periodically repressurized with oxygen from Progress 27P.)

FE-2 Tani performed his daily status check on the BCAT-3 (Binary Colloidal Alloy Test-3) science payload, running by itself in Node-2 since 12/13/07 (briefly interrupted for EVA-13 photo support). (The status check, conducted on the last image taken by the DCS 760 digital still camera which is controlled by EarthKAM software on an A31p laptop, is to verify proper image focus and camera alignment. The SSC (Station Support Computer) is taking photography of the phase separation occurring in the BCAT Sample 3, with the photo flash going off every half hour. Special uplink to Dan this morning: 'The BCAT team downloaded the new images and reports that they look fabulous. The team is very excited and looking forward to pressing ahead and making more progress in coming days.')

FE-1 Malenchenko meanwhile conducted the routine servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. (Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers. Weekly SOZh reports (on Sundays) to TsUP/Moscow deal with number & dates of water and urine containers, counter readings of water consumption & urine collection, and total operating time of the POTOK air filtration system.)

Working off his 'time permitting' discretionary task list, Yuri conducted his fifth run of the Russian DZZ-2 "Diatomeya" ocean observations program, using the NIKON-F5 still camera with 80-200 mm Nikkor zoom lens and the SONY PD-150O camcorder to record color contrasts on water surface and irregular cloud cover pattern in the World Ocean target areas. (Uplinked target zones in the Atlantic Ocean were the Great Banks, the commercial-production area of the Western Sahara, the US coastal area of the Agulhas subsea table (RSA), the Gulf of Mexico with the Caribbean Sea and the Venezuelan coast.)

A second job item on the FE-1's discretionary list for today was another KPT-3 session to make observations and take aerial KPT-3 photography of environmental conditions for Russia's Environmental Safety Agency (ECON) using the Nikon D2X digital camera with SIGMA 300-800mm telephoto lens. (Targets today were contamination areas in the Seattle area and in the Pacific Ocean.)

A new entry added to the discretionary US 'job jar' task list for Peggy and Yuri is 'ghosting' an updated personal image with their preferred material on an 60GB hard disk in an A31p laptop UltraBay Adapter reserved as their CPSD (Crew Personal Support Disk), a 45-min task for each.

Also on Whitson's voluntary task lit, at her convenience, is an audit of rack locations, using the IMS (Inventory Management System). (The audit function, introduced with the implementation of IMS software version 2.0, allows the crewmember to set up audits of bags, kits, containers and stowage locations on the laptop.)

Another newly added item on the 'job jar' task list is for FE-02 Tani to continue his crew departure preparations.

The crewmembers performed their regular 2.5-hr physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2), TVIS treadmill (CDR , FE-1), RED (CDR, FE-2) and VELO bike with bungee cord load trainer (FE-1).

Tani and Whitson had their weekly PFCs (Private Family Conferences) via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on the SSC-9 laptop), Dan at ~9:20am EST, Peggy at ~10:55am.

No CEO photo targets uplinked for today.

More... - Chronology...

Home - Browse - Contact
© / Conditions for Use